OpenStudy (anonymous):

The perimeter of a rectangle is 96 ft. The ratio of its length to its width is 7 : 5. What are the dimensions of the rectangle? A. 30 ft by 18 ft B. 28 ft by 20 ft C. 24 ft by 24 ft D. 22 ft by 26 ft

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@Jonask

OpenStudy (anonymous):

Okay, since they give you the ratio and its multiple choice, all you have to do is find which numbers that are given as answers, are divisible by 7 and 5. So, which is the answer?

OpenStudy (anonymous):

Divisible is a number that is capable of being divided by another number without a remainder: "24 is divisible by 4".

OpenStudy (anonymous):

You could not be more correct:) Good job!

OpenStudy (anonymous):

\[2(l+b)=96\] \[l/b=7/5\]------------\[5l=7b\] solve for a and b